a woodland jumping mouse hops along a parabolic path given by y=-0.2xsquared +1.3x where x is the mouses horazontal position(in feet) and y is the corasponding hight (in feet). can the mouse jump over a fence that is 3 feet high? explain
Answer by stanbon(75887)

a woodland jumping mouse hops along a parabolic path given by
y=-0.2xsquared +1.3x
where x is the mouse's horizontal position(in feet) and y is the corasponding height (in feet).
can the mouse jump over a fence that is 3 feet high? explain
------
Maximum height occurs at x = -b/2a = -1.3/(2*-0.2) = 3.25 ft
----
The mouse reaches maximum height at 3.25 ft from its starting point.
---
Its height at that position is
f(3.25) = -0.2(3.25)^2+1.3(3.25) = 2.1125 ft
-------------------------------
Mousy cannot get over the fence.
-------------------------------------
Cheers,
Stan H.